OpenStudy (anonymous):

please help! :)) 2. Solve for y: y^2 + y + 1 = 0 using the quadratic formula I don't know how to do this so bare with me:)

OpenStudy (mertsj):

Do you know the formula?

OpenStudy (anonymous):

do you allow complex solutions? because the answer will not be real

OpenStudy (anonymous):

y = [-b ± √(b^2 - 4ac)] / 2a

OpenStudy (anonymous):

yup use \(a=1,b=1,c=1\)

OpenStudy (anonymous):

and @satellite73 I have no clue

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

the number under the radical will be \(1-4=-3\) so the solutions are complex

OpenStudy (anonymous):

write them directly in to your formula get \[\frac{-1\pm\sqrt{1-4}}{2}\] as a first step

OpenStudy (anonymous):

then since \(1-4=-3\) you get \[\frac{-1\pm\sqrt{-3}}{2}\]

OpenStudy (anonymous):

okay so you basically cleaned it up

OpenStudy (anonymous):

if you are working with complex numbers you can rewrite this as \[\frac{-1\pm\sqrt{3}i}{2}\] and if you are not, you say "no solution"

OpenStudy (anonymous):

yeah the only clean up is \(1-4=-3\) that is really all the computation necessary

OpenStudy (pradius):

\[Ax ^{2}+Bx+C\] Quadratic formula: \[x= \frac{-B+-\sqrt{B ^{2}-4AC} }{ 2A }\]

OpenStudy (anonymous):

but i'm a tad bit confused?

OpenStudy (anonymous):

i'm trying to find y

OpenStudy (anonymous):

yeah, that is \(y\)

OpenStudy (anonymous):

actually you get two answer, the plus and the minus i.e. either \[y=\frac{-1+\sqrt{-3}}{2}\] or \[y=\frac{-1-\sqrt{-3}}{2}\] those are your two solutions

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

okay so i basically understand except how you got the number under the radical

OpenStudy (anonymous):

the number under the radical is \(b^2-4ac\) in your case \(a=1\) and \(b=1\) and even \(c=1\) so literally you get \(1^2-4\times 1\times 1\) under the radical

OpenStudy (anonymous):

but this is really \(1-4=-3\)

OpenStudy (anonymous):

if you got another one we can try that and see how it works in a different example

OpenStudy (anonymous):

I'm sorta getting the concept. I'm gonna write this down and look at it more to understand it and that was the only quadratic problem so I don't have another example.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thank you so much though:)

OpenStudy (anonymous):

yw